OpenStudy (anonymous):

an express bus travels 320 miles in the same amount of time that a car travels 280 miles. The rate of the car is 8 mph less than the rate of the bus. Find the rate of the bus.... please help!!!

OpenStudy (anonymous):

Rate x Time = Distance You know that time is the same for both vehicles. You want to change the formula in terms of Time. So: Time= Distance/Rate. Then, set the two equations equal to each other and solve.

OpenStudy (anonymous):

\[\frac{ 320 }{ x+8 }=\frac{ 280 }{ x-8 }?\]

OpenStudy (anonymous):

Sorry for the delay. You made a common mistake."The rate of the car is 8 mph less than the rate of the bus." Let's say that the rate of the bus is x. Then we could substitute the part about the rate of the bus as "The rate of the car is 8 mph less than x". Can you catch what you did wrong?

OpenStudy (anonymous):

\[\frac{ 320 }{ x }\]

OpenStudy (anonymous):

Exactly. so 320/x=280/(x-8)

OpenStudy (anonymous):

Solve for x and tell me just in case.

OpenStudy (anonymous):

x=64

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

thank you